OpenStudy (anonymous):

Find the vertex -16x^2 + 20,504x

OpenStudy (anonymous):

hi friend

OpenStudy (anonymous):

can u hurry up

OpenStudy (anonymous):

its like rabies

OpenStudy (anonymous):

if by vertex you mean the stationary point of that curve: \[y = -16x^2 + 20504x\]\[\frac{dy}{dx}= -32x +20504\] set the second equation equal to zero to find the point where the gradient of tangent = 0 \[ -32x +20504 =0\]\[x = \frac{20504}{32} = 640.75\] this is the x coordinate, substitute into the first equation to get the y coordinate.

OpenStudy (anonymous):

Correct method finding the derivative, but for a quadratic function, you can use the relationship x=-b/2a to find the axis of symmetry (the x coordinate of the vertex).

OpenStudy (anonymous):

yes, although i would argue knowing where that relationship comes from is more important than putting numbers in a formula

OpenStudy (anonymous):

You are correct, but for an algebra student, the calculus might not be very accessible.

OpenStudy (anonymous):

fair point

OpenStudy (anonymous):

is there a geometric derivation of -b/2a ? im interested

OpenStudy (anonymous):

Comes from the process of completing the square. The first term is -b/2a plus or minus something. That puts the axis of symmetry there, midway between the two roots.